OpenStudy (anonymous):

p(less than 3) on a spinner that is divided into 10 sectors labeled 1 through 5 ha are paced into two different sectors what is the probbility of it

OpenStudy (anonymous):

Do I understand it properly that there are two sectorst labeled 1, two are labeled 2 .... and two are labeled 5? And that two things are each placed in one of the sectors and the quaestion asks what the probability of the sum of thows two sectors is less than 3?

OpenStudy (anonymous):

Well in that case there is only one way to have \(P(x<3)\) and that is when the first pawn is on a sector labelled one and after that the other one is also placed on the other labelled 1 THe chance that the first pawn is on a sector labelled one is \(\frac{ 2 }{ 10 } = \frac{ 1 }{ 5 }\) Now 9 sectors are left and only one is labelled 1, so the chance of the second pawn to be on the one labelled 1 is onle \(\frac{ 1 }{ 9 } \) The chance of this both combined will be the product of those chances is \[ P(x<3)=\frac{ 1 }{ 5 } \times \frac{ 1 }{ 9 } = \frac{ 1 }{ 45 }\]